Break-Even Sales and Cost-Volume-Profit Graph

For the coming year, Bernardino Company anticipates a unit selling price of $140, a unit variable cost of $70, and fixed costs of $735,000.

Instructions:

1. Compute the anticipated break-even sales in units.
_________________ units

2. Compute the sales (units) required to realize operating income of $322,000.
_________________ units

3. Construct a cost-volume-profit graph on paper, assuming maximum sales of 21,000 units within the relevant range. From your chart, indicate whether each of the following sales levels would produce a profit, a loss, or break-even.

$2,058,000 (Break Even, Loss, or Profit)
$1,834,000 (Break Even, Loss, or Profit)
$1,470,000 (Break Even, Loss, or Profit)
$1,106,000 (Break Even, Loss, or Profit)
$882,000 (Break Even, Loss, or Profit)

4. Determine the probable operating income (loss) if sales total 16,800 units. If required, use the minus sign to indicate a loss.
$ ___________________ ( Profit or Loss)
Answer #1

1) Break even unit = Fixed cost/Contribution margin per unit = 735000/(140-70) = 10500 Units

2) Required sales unit = (735000+322000)/70 = 15100 Units

3)

$2,058,000 Profit
$1,834,000 Profit
$1,470,000 Break even
$1,106,000 Loss
$882,000 Loss

4) Net income (loss) = 16800*70- 735000 = 441000